OpenStudy (anonymous):

A compound containing only carbon and hydrogen has a carbon-to-hydrogen mass ratio of 11.89. Which carbon-to-hydrogen mass ratio is possible for another compound composed of only carbon and hydrogen? I'm not quite sure how to go about solving this. Any help would be appreciated.

OpenStudy (aaronq):

Are there options? you're basically using a ratio of the molar masses and the number of atoms, but this will vary depending on the type of bonding. \(\sf ratio=\dfrac{(molar~mass~of~C)*(number ~of~C~atoms)}{(molar~mass~of~H)*(number~of~H~atoms )}\) Pluggin in the molar masses reduces this to, \(\sf ratio=\dfrac{ {12*(number ~of~C~atoms)}}{{1*(number~of~H~atoms )}}=12\dfrac{ {(number ~of~C~atoms)}}{{(number~of~H~atoms )}}\) Plug in the number of atoms for the different types of hydrocarbons (e.g. alkanes, alkenes)

OpenStudy (anonymous):

You're amazing, thank you so much. This question had me stumped!